OpenStudy (anonymous):

Write a function with the following characteristics: A vertical asymptote at x=-1 An oblique asymptote at y=x+2

OpenStudy (freckles):

for there to be a oblique asymptote the degree of numerator needs to be one higher than the degree of the denominator

OpenStudy (freckles):

you also want the bottom to contain the factor (x+1) since you want a vertical asymptote at x=-1

OpenStudy (freckles):

so let's assume the bottom is x+1 then the top being one degree higher would be of the form ax^2+bx+c

OpenStudy (freckles):

\[\frac{ax^2+bx+c}{x+1}=x+2\]

OpenStudy (freckles):

find a,b, and c

OpenStudy (freckles):

and actually we still have a problem

OpenStudy (anonymous):

Okay wow that was a lot thank you but I don't know what a, b, and c would be given the information that I have

OpenStudy (freckles):

that would give us a hole at x=-1 not a vertical asymptote

OpenStudy (anonymous):

Okay so what would this problem look like?

OpenStudy (anonymous):

@freckles

OpenStudy (freckles):

\[\frac{ax^2+bx+c}{x+1}=x+2+\frac{R}{x+1}\] there assuming R is not 0 that means x+1 is not a factor of ax^2+bx+c problem gotten rid of

OpenStudy (anonymous):

okay so that would be the answer then? @freckles

OpenStudy (freckles):

no you have to solve for a,b, and c

OpenStudy (anonymous):

I still don't know what a, b, and c are.

OpenStudy (freckles):

have you tried multiplying both sides by (x+1)

OpenStudy (freckles):

and expanding and then comparing both sides that is one way

OpenStudy (freckles):

can you show me how far you gotten with this?

OpenStudy (anonymous):

Well Im basically don't have anything so I can't show you how far I am

OpenStudy (freckles):

so you haven't tried doing any of the things I asked you to?

OpenStudy (anonymous):

Im in the middle of doing it just its not coming out right so I'm having to redo it again

OpenStudy (freckles):

can you show me what you have after multiplying both sides of that equation above the equation \[\frac{ax^2+bx+c}{x+1}=x+2+\frac{R}{x+1}\] by (x+1) this was the first step I asked you to do above

OpenStudy (freckles):

\[\frac{P(x)}{x+1}=Q(x)+\frac{R}{x+1} \\ \text{ multiply both sides by } x+1 \\ P(x)=Q(x)(x+1)+R\] then expand the Q(x)(x+1) part in your problem Q(x) being (x+2) here

OpenStudy (anonymous):

Okay so I don't know what you are doing. I did not learn this in my unit so I think that is why I am so confused at what you are saying.

OpenStudy (freckles):

oh you doing know how to multiply (x+1) on both sides?

OpenStudy (freckles):

\[(x+1) \cdot \frac{ax^2+bx+c}{x+1}=(x+1) [x+2+\frac{R}{x+1}] \\ (ax^2+bx+c) \cdot \frac{x+1}{x+1}=(x+1)(x+2)+(x+1) \cdot \frac{R}{x+1} \text{ by distributive property } \\ (ax^2+bx+c) \cdot 1 =(x+1)(x+2) +R \cdot \frac{x+1}{x+1} \\ ax^2+bx+c=(x+1)(x+2)+R \cdot 1 \\ ax^2+bx+c=(x+1)(x+2)+R\]

OpenStudy (freckles):

expand the (x+1)(x+2)

OpenStudy (freckles):

and then compare both sides to find what a,b, and c are

OpenStudy (freckles):

you will see there are infinitely many choices for c

OpenStudy (freckles):

have you expanded (x+1)(x+2)?

OpenStudy (freckles):

to multiply (x+1)(x+2) just use distributive property twice (x+1)(x+2) =x(x+2)+1(x+2) =....

OpenStudy (anonymous):

x^2+3x+2

OpenStudy (freckles):

great so you have \[ax^2+bx+c=x^2+3x+2+R\]

OpenStudy (freckles):

compare both sides a=? b=? c=?

OpenStudy (anonymous):

a= 1x^2 b=3x c=2

OpenStudy (freckles):

well no that means you have x^2*x^2+3x*x+2=x^2+3x+2+R which is not true

OpenStudy (freckles):

in order for this equation to hold you need to pick a,b,c such that you have the same about of x^2 on both sides the same amount of x on both sides and the constant on both sides

OpenStudy (freckles):

for example ax^2=x^2 when a=1

OpenStudy (anonymous):

a=1 b=3 c=2

OpenStudy (freckles):

a and b are correct but for c you ignored the R

OpenStudy (anonymous):

c=2+R

OpenStudy (freckles):

\[ax^2+bx+c=x^2+3x+2+R \\ \implies \\ a=1 \\ b=3 \\ c=2+R\] where we said R couldn't be 0

OpenStudy (freckles):

so as long as you choose R not to be 0 you are fine in your choice for c

OpenStudy (anonymous):

okay so what is the final answer then?

OpenStudy (freckles):

have you chose the value of R you want to use yet?

OpenStudy (freckles):

you have infinitely many numbers to choose from except 0 as I said

OpenStudy (anonymous):

oh okay well I got it thank you for your help

OpenStudy (freckles):

\[\frac{ax^2+bx+c}{x+1} \text{ has } \text{ va } x=-1 \text{ and } \text{ oba } y=x+2 \\ \text{ if you \choose } a=1,b=3, c=2+R , R \neq 0\] so R=1 would work R=12 or R=-12 R=1/2 whatever just you can't choose R to be 0